# A novel multipoint stress control volume method for linear elasticity on quadrilateral grids

### v1

@@ 4.1 Some elementary properties
$\sum_{e\in\mathcal{F}_{pr}^{\frac{1}{2}}}(\underline{w}_{h}\bm{n},\left\llbracket\bm{v}_{h}\right\rrbracket)_{e}+(\text{as}(\underline{w}_{h}),\eta_{h})=\sum_{e\in\mathcal{F}_{pr}^{\frac{1}{2}}}h_{e}^{-1}\|\left\llbracket\bm{v}_{h}\right\rrbracket\|_{L^{2}(e)}^{2}+\|\eta_{h}\|_{L^{2}(\Omega)}^{2}.$
This, (4.17) and (4.20) complete the proof.
@@
The methodologies used in the proof Lemma 4.2 hold for both Method 1 and Method 2, and there is only subtle difference in deriving (4.19).
@@
(existence and uniqueness). There exists a unique solution to (3.2)-(3.3). In addition, there exists a positive constant $C$ independent of $\lambda$ such that the following estimate holds
@@
Taking $\underline{w}_{h}=\underline{I}$ in (3.2), we have
$(\mathcal{A}\underline{\sigma}_{h},\underline{I})=\frac{1}{2(\mu+\lambda)}\int_{\Omega}\mathrm{tr}(\underline{\sigma}_{h})=0.$
Thus, $\mathrm{tr}(\underline{\sigma}_{h})\in L^{2}_{0}(\Omega)$. Then, there exists $\bm{\psi}\in\bm{H}^{1}_{0}(\Omega)$ (cf. [21]) such that
$\nabla\cdot\bm{\psi}=\mathrm{tr}(\underline{\sigma}_{h}),\quad\|\bm{\psi}\|_{H^{1}(\Omega)}\leq C\|\mathrm{tr}(\underline{\sigma}_{h})\|_{L^{2}(\Omega)}.$ (4.23)
Then we have
$\|\mathrm{tr}(\underline{\sigma}_{h})\|_{L^{2}(\Omega)}^{2}=(\mathrm{tr}(\underline{\sigma}_{h}),\nabla\cdot\bm{\psi})=2(\underline{\sigma}_{h}-\mathrm{dev}\underline{\sigma}_{h},\nabla\bm{\psi}).$
An application of integration by parts and the Cauchy-Schwarz inequality gives
$(\underline{\sigma}_{h},\nabla\bm{\psi})$ $=\sum_{e\in\mathcal{F}_{dl}}(\left\llbracket\underline{\sigma}_{h}\bm{n}\right\rrbracket,\bm{\psi})_{e}=\sum_{e\in\mathcal{F}_{dl}}(\left\llbracket\underline{\sigma}_{h}\bm{n}\right\rrbracket,\mathcal{J}_{h}\bm{\psi})_{e}=(\bm{f},\mathcal{J}_{h}\bm{\psi})\leq C\|\bm{f}\|_{L^{2}(\Omega)}\|\mathcal{J}_{h}\bm{\psi}\|_{L^{2}(\Omega)}.$
This and (4.23) yield (4.21). Since (3.2)-(3.3) is a square linear system, the uniqueness implies the existence. Setting $\bm{f}=\bm{0}$ in (4.21) implies $\underline{\sigma}_{h}=\underline{0},\bm{u}_{h}=0$ and $\gamma_{h}=0$.
@@
4.2 Quasi-optimal approximation
Inspired by [30], we introduce the following moment-preserving operator that is crucial for establishing the quasi-optimal approximation without requiring additional regularity assumption.
Lemma 4.4.
(moment-preserving operator). Let $E:\bm{U}_{h}\rightarrow\bm{H}^{1}_{0}(\Omega)$ satisfy
$\int_{e}(E\bm{v})\bm{q}=\int_{e}\bm{v}\bm{q}\quad\forall\bm{q}\in\bm{P}_{0}(e),\quad\forall e\in\mathcal{F}_{h}.$
In addition, for all $\bm{v}\in\bm{U}_{h}$, we have
$\|\nabla(E\bm{v})\|_{L^{2}(\Omega)}\leq C\|\bm{v}\|_{1,h}.$
Proof.
The proof can be found in [30], which is omitted here for the sake of simplicity. ∎
We let $\text{dev}(\underline{w})=\underline{w}-\frac{1}{d}\mathrm{tr}(\underline{w})\underline{I}$. In view of [17, Lemma 3.7(b)], we have
Lemma 4.5.
Let $\underline{\Sigma}_{0}:=\{\underline{\tau}\in\underline{L}^{2}(\Omega):\int_{\Omega}\mathrm{tr}(\underline{\tau})=0\}$. For $\underline{w}\in\underline{\Sigma}_{0}$, it holds
$\|\mathrm{tr}(\underline{w})\|_{L^{2}(\Omega)}$ $\leq C\Big{(}\|\textnormal{dev}(\underline{w})\|_{L^{2}(\Omega)}+\|\mathrm{div}\underline{w}\|_{-1}\Big{)},$ (4.24)
$\|\underline{w}\|_{L^{2}(\Omega)}$ $\leq C\Big{(}\mu^{1/2}\|\underline{w}\|_{\mathcal{A}}+\|\mathrm{div}\underline{w}\|_{-1}\Big{)}.$ (4.25)
Lemma 4.6.
($\|\mathrm{div}(\underline{\sigma}-\underline{\sigma}_{h})\|_{-1}$). The following error estimate holds
$\|\mathrm{div}(\underline{\sigma}-\underline{\sigma}_{h})\|_{-1}\leq C\Big{(}\sum_{M\in\mathcal{T}_{M}}h_{M}^{2}\|\bm{f}\|_{L^{2}(M)}^{2}\Big{)}^{\frac{1}{2}}.$
Proof.
For $\bm{v}\in\bm{H}^{1}_{0}(\Omega)$, we can infer from (4.3) and (4.4) that
$\begin{split}(\underline{\sigma}-\underline{\sigma}_{h},\nabla\bm{v})&=(\bm{f},\bm{v})-\sum_{e\in\mathcal{F}_{dl}}(\left\llbracket\underline{\sigma}_{h}\bm{n}\right\rrbracket,\bm{v})_{e}=(\bm{f},\bm{v})-\sum_{e\in\mathcal{F}_{dl}}(\left\llbracket\underline{\sigma}_{h}\bm{n}\right\rrbracket,\mathcal{J}_{h}\bm{v})_{e}\\ &=(\bm{f},\bm{v}-\mathcal{J}_{h}\bm{v})\leq C\Big{(}\sum_{M\in\mathcal{T}_{M}}h_{M}^{2}\|\bm{f}\|_{L^{2}(M)}^{2}\Big{)}^{\frac{1}{2}}\|\nabla\bm{v}\|_{L^{2}(\Omega)},\end{split}$ (4.26)
which combined with the definition of the negative norm completes the proof.
Theorem 4.2.
(quasi-optimal approximation for $\|\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}$). Let $(\underline{\sigma},\bm{u},\gamma)$ be the exact solution and let $(\underline{\sigma}_{h},\bm{u}_{h},\gamma_{h})\in\underline{\Sigma}_{h}\times\bm{U}_{h}\times\Gamma_{h}$ be the numerical solution of (3.2)-(3.4) obtained by either Method 1 or Method 2. It holds
$\|\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}\leq C\Big{(}\inf_{\widehat{\underline{\sigma}}_{h}\in\underline{\Sigma}_{h}}\|\widehat{\sigma}_{h}-\underline{\sigma}\|_{L^{2}(\Omega)}+\inf_{\widehat{\gamma}_{h}\in\Gamma_{h}}\|\gamma-\widehat{\gamma}_{h}\|_{L^{2}(\Omega)}+\Big{(}\sum_{M\in\mathcal{T}_{M}}h_{M}^{2}\|\bm{f}\|_{L^{2}(M)}^{2}\Big{)}^{\frac{1}{2}}\Big{)}.$
Proof.
Let $\widehat{\underline{\sigma}}_{h}\in\underline{\Sigma}_{h}$ and $\widehat{\gamma}_{h}\in\Gamma_{h}$ be arbitrary functions. We let $\underline{e}_{\sigma}=\widehat{\underline{\sigma}}_{h}-\underline{\sigma}_{h}$, $\bm{e}_{u}=\mathcal{J}_{h}\bm{u}-\bm{u}_{h}$ and $e_{\gamma}=\widehat{\gamma}_{h}-\gamma_{h}$. Then
$\|\underline{\sigma}-\underline{\sigma}_{h}\|_{\mathcal{A}}^{2}=(\mathcal{A}(\underline{\sigma}-\underline{\sigma}_{h}),\underline{e}_{\sigma})+(\mathcal{A}(\underline{\sigma}-\underline{\sigma}_{h}),\underline{\sigma}-\widehat{\underline{\sigma}}_{h}).$ (4.27)
An application of integration by parts and (3.2) leads to
$\begin{split}(\mathcal{A}(\underline{\sigma}-\underline{\sigma}_{h}),\underline{e}_{\sigma})&=(\nabla\bm{u},\underline{e}_{\sigma})-\sum_{e\in\mathcal{F}_{dl}}(\bm{u}_{h},\left\llbracket\underline{e}_{\sigma}\bm{n}\right\rrbracket)_{e}-(\text{as}(\underline{e}_{\sigma}),\gamma)+(\text{as}(\underline{e}_{\sigma}),\gamma_{h})\\ &=\sum_{e\in\mathcal{F}_{dl}}(\bm{u},\left\llbracket\underline{e}_{\sigma}\bm{n}\right\rrbracket)_{e}-\sum_{e\in\mathcal{F}_{dl}}(\bm{u}_{h},\left\llbracket\underline{e}_{\sigma}\bm{n}\right\rrbracket)_{e}-(\text{as}(\underline{e}_{\sigma}),\gamma-\gamma_{h})\\ &=\sum_{e\in\mathcal{F}_{dl}}(\mathcal{J}_{h}\bm{u}-\bm{u}_{h},\left\llbracket\underline{e}_{\sigma}\bm{n}\right\rrbracket)_{e}-(\text{as}(\underline{e}_{\sigma}),\gamma-\widehat{\gamma}_{h})-(\text{as}(\underline{e}_{\sigma}),\widehat{\gamma}_{h}-\gamma_{h}).\end{split}$ (4.28)
Owing to $(\text{as}(\underline{\sigma}-\underline{\sigma}_{h}),\xi_{h})=0,\forall\xi_{h}\in\Gamma_{h}$, we have
$\begin{split}(\text{as}(\underline{e}_{\sigma}),\widehat{\gamma}_{h}-\gamma_{h})&=(\text{as}(\underline{e}_{\sigma}),\widehat{\gamma}_{h}-\gamma_{h})=(\text{as}(\widehat{\underline{\sigma}}_{h}-\underline{\sigma}),\widehat{\gamma}_{h}-\gamma_{h})+(\text{as}(\underline{\sigma}-\underline{\sigma}_{h}),\widehat{\gamma}_{h}-\gamma_{h}).\end{split}$ (4.29)
On the other hand, an application of integration by parts gives
$(\underline{e}_{\sigma},\nabla(E\bm{e}_{u}))=\sum_{e\in\mathcal{F}_{dl}}(\left\llbracket\underline{e}_{\sigma}\bm{n}\right\rrbracket,E\bm{e}_{u})_{e}=\sum_{e\in\mathcal{F}_{dl}}(\left\llbracket\underline{e}_{\sigma}\bm{n}\right\rrbracket,\bm{e}_{u})_{e}.$
This, (4.28) and (4.29) yield
$\begin{split}(\mathcal{A}(\underline{\sigma}-\underline{\sigma}_{h}),\underline{e}_{\sigma})&=(\underline{e}_{\sigma},\nabla(E\bm{e}_{u}))-(\text{as}(\underline{e}_{\sigma}),\gamma-\widehat{\gamma}_{h})-(\text{as}(\underline{e}_{\sigma}),\widehat{\gamma}_{h}-\gamma_{h})\\ &=(\widehat{\sigma}_{h}-\underline{\sigma},\nabla(E\bm{e}_{u}))+(\underline{\sigma}-\underline{\sigma}_{h},\nabla(E\bm{e}_{u}))-(\text{as}(\underline{e}_{\sigma}),\gamma-\widehat{\gamma}_{h})-(\text{as}(\widehat{\sigma}_{h}-\underline{\sigma}),\widehat{\gamma}_{h}-\gamma_{h}).\end{split}$ (4.30)
The Cauchy-Schwarz inequality and Lemma 4.4 give
$\begin{split}(\widehat{\sigma}_{h}-\underline{\sigma},\nabla(E\bm{e}_{u}))&\leq\|\widehat{\sigma}_{h}-\underline{\sigma}\|_{L^{2}(\Omega)}\|\nabla(E\bm{e}_{u})\|_{L^{2}(\Omega)}\leq C\|\widehat{\sigma}_{h}-\underline{\sigma}\|_{L^{2}(\Omega)}\|\bm{e}_{u}\|_{1,h}.\end{split}$ (4.31)
The Poincaré-Friedrichs inequality yields
$\begin{split}(\underline{\sigma}-\underline{\sigma}_{h},\nabla(E\bm{e}_{u}))&=(\bm{f},E\bm{e}_{u})-\sum_{e\in\mathcal{F}_{dl}}(\left\llbracket\underline{\sigma}_{h}\bm{n}\right\rrbracket,E\bm{e}_{u})_{e}=(\bm{f},E\bm{e}_{u})-\sum_{e\in\mathcal{F}_{dl}}(\left\llbracket\underline{\sigma}_{h}\bm{n}\right\rrbracket,\bm{e}_{u})_{e}\\ &=(\bm{f},E\bm{e}_{u}-\bm{e}_{u})\leq C\Big{(}\sum_{M\in\mathcal{T}_{M}}h_{M}^{2}\|\bm{f}\|_{L^{2}(M)}^{2}\Big{)}^{\frac{1}{2}}\|\nabla(E\bm{e}_{u})\|_{L^{2}(\Omega)}\\ &\leq C\Big{(}\sum_{M\in\mathcal{T}_{M}}h_{M}^{2}\|\bm{f}\|_{L^{2}(M)}^{2}\Big{)}^{\frac{1}{2}}\|\bm{e}_{u}\|_{1,h}.\end{split}$ (4.32)
In view of (4.28) and (4.3), we have
$(\mathcal{A}(\underline{\sigma}-\underline{\sigma}_{h},\underline{w})+\sum_{e\in\mathcal{F}_{dl}}(\mathcal{J}_{h}\bm{u}-\bm{u}_{h},\left\llbracket\underline{w}\bm{n}\right\rrbracket)_{e}+(\gamma-\gamma_{h},\text{as}(\underline{w}))$ $=0\quad\forall\underline{w}\in\underline{\Sigma}_{h}.$
This and the inf-sup condition (4.2) yield
$\begin{split}\|(\bm{e}_{u},e_{\gamma})\|_{h}&\leq C\sup_{\underline{0}\neq\underline{w}_{h}\in\underline{\Sigma}_{h}}\frac{\sum_{e\in\mathcal{F}_{pr}^{\frac{1}{2}}}(\underline{w}_{h}\bm{n},\left\llbracket\bm{e}_{u}\right\rrbracket)_{e}+(\text{as}(\underline{w}_{h}),e_{\gamma})}{\|\underline{w}_{h}\|_{L^{2}(\Omega)}}\\ &\leq C\Big{(}\|\mathcal{A}(\underline{\sigma}-\underline{\sigma}_{h})\|_{L^{2}(\Omega)}+\|\gamma-\widehat{\gamma}_{h}\|_{L^{2}(\Omega)}\Big{)}.\end{split}$ (4.33)
This, (4.27) , (4.30), (4.32) and (4.31) imply
$\|\underline{\sigma}-\underline{\sigma}_{h}\|_{\mathcal{A}}\leq C\Big{(}\|\widehat{\sigma}_{h}-\underline{\sigma}\|_{L^{2}(\Omega)}+\|\gamma-\widehat{\gamma}_{h}\|_{L^{2}(\Omega)}+\Big{(}\sum_{M\in\mathcal{T}_{M}}h_{M}^{2}\|\bm{f}\|_{L^{2}(M)}^{2}\Big{)}^{\frac{1}{2}}\Big{)}.$
In view of $\mathcal{A}\underline{\sigma}=\nabla\bm{u}-\gamma\underline{\delta}$, we can infer from (3.2) that
$(\mathcal{A}(\underline{\sigma}-\underline{\sigma}_{h},\underline{w})+\sum_{e\in\mathcal{F}_{dl}}(\bm{u}-\bm{u}_{h},\left\llbracket\underline{w}\bm{n}\right\rrbracket)_{e}+(\gamma-\gamma_{h},\text{as}(\underline{w}))$ $=0\quad\forall\underline{w}\in\underline{\Sigma}_{h}.$ (4.34)
[…]
$\|\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}\leq C\Big{(}\|\widehat{\sigma}_{h}-\underline{\sigma}\|_{L^{2}(\Omega)}+\|\gamma-\widehat{\gamma}_{h}\|_{L^{2}(\Omega)}+\Big{(}\sum_{M\in\mathcal{T}_{M}}h_{M}^{2}\|\bm{f}\|_{L^{2}(M)}^{2}\Big{)}^{\frac{1}{2}}\Big{)}.$
Thus, the proof is completed. ∎
In view of (4.33) and Theorem 4.2, we can infer the following corollary.
Corollary 4.1.
(quasi-optimal approximation for $\|(\bm{e}_{u},e_{\gamma})\|_{h}$). Let $(\underline{\sigma},\bm{u},\gamma)$ be the exact solution and let $(\underline{\sigma}_{h},\bm{u}_{h},\gamma_{h})\in\underline{\Sigma}_{h}\times\bm{U}_{h}\times\Gamma_{h}$ be the numerical solution of (3.2)-(3.4) obtained by either Method 1 or Method 2. Then, it holds
$\|(\bm{e}_{u},e_{\gamma})\|_{h}\leq C\Big{(}\inf_{\widehat{\sigma}_{h}\in\underline{\Sigma}_{h}}\|\widehat{\sigma}_{h}-\underline{\sigma}\|_{L^{2}(\Omega)}+\inf_{\widehat{\gamma}_{h}\in\Gamma_{h}}\|\gamma-\widehat{\gamma}_{h}\|_{L^{2}(\Omega)}+\Big{(}\sum_{M\in\mathcal{T}_{M}}h_{M}^{2}\|\bm{f}\|_{L^{2}(M)}^{2}\Big{)}^{\frac{1}{2}}\Big{)}.$
If $\underline{\sigma}\in\underline{H}^{1}(\Omega)$ and $\gamma\in H^{1}(\Omega)$, then the following convergence holds
$\|\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}$ $\leq C\Big{(}h|\underline{\sigma}|_{H^{1}(\Omega)}+h|\gamma|_{H^{1}(\Omega)}+\Big{(}\sum_{M\in\mathcal{T}_{M}}h_{M}^{2}\|\bm{f}\|_{L^{2}(M)}^{2}\Big{)}^{\frac{1}{2}}\Big{)}.$ (4.35)
Moreover, the following superconvergence holds
$\|\bm{e}_{u}\|_{1,h}\leq C\Big{(}h|\underline{\sigma}|_{H^{1}(\Omega)}+h|\gamma|_{H^{1}(\Omega)}+\Big{(}\sum_{M\in\mathcal{T}_{M}}h_{M}^{2}\|\bm{f}\|_{L^{2}(M)}^{2}\Big{)}^{\frac{1}{2}}\Big{)}.$
Lemma 4.7.
(inf-sup condition). There exists a positive constant $C$ such that
$\sup_{\bm{0}\neq\underline{w}\in\underline{\Sigma}_{h}}\frac{(\bm{v},\mathrm{div}\underline{\Pi}_{h}\underline{w})}{\|\underline{\Pi}_{h}\underline{w}\|_{\mathrm{div}}}\geq C\|\bm{v}\|_{L^{2}(\Omega)},\quad\sup_{\bm{0}\neq\underline{w}\in\underline{\Sigma}_{h}}\frac{(\bm{v},\mathrm{div}\underline{\Pi}_{h}\underline{w})}{\|\underline{w}\|_{\Sigma,h}}\geq C\|\bm{v}\|_{L^{2}(\Omega)}\quad\forall\bm{v}\in\bm{U}_{h},$
where
$\|\underline{w}\|_{\Sigma,h}^{2}=\sum_{e\in\mathcal{F}_{dl}}h_{e}^{-1}\|\left\llbracket\underline{w}\right\rrbracket\|_{L^{2}(e)}^{2}+\sum_{e\in\mathcal{F}_{pr}^{\frac{1}{2}}}h_{e}^{-1}\|\left\llbracket\underline{w}\right\rrbracket\|_{L^{2}(e)}^{2}.$
Proof.
$\|\underline{I}_{h}^{e}\underline{\theta}\|_{\Sigma,h}\leq Ch^{-1}\|\underline{\theta}-\underline{I}_{h}^{e}\underline{\theta}\|_{L^{2}(\Omega)}+\|\nabla\underline{\theta}\|_{L^{2}(\Omega)}\leq C\|\underline{\theta}\|_{H^{1}(\Omega)}.$ (4.37)
We can infer from (4.36) and (4.5) that
Then we have
$\begin{split}(\bm{v},\mathrm{div}\underline{\Pi}_{h}\underline{\theta})&=(\bm{v},\mathrm{div}\underline{\Pi}_{h}\underline{I}_{h}^{e}\underline{\theta})=\sum_{M\in\mathcal{T}_{M}}(\bm{v},\underline{\Pi}_{h}\underline{I}_{h}^{e}\underline{\theta}\bm{n})_{\partial M}\\ &=\sum_{M\in\mathcal{T}_{M}}(\bm{v},\underline{I}_{h}^{e}\underline{\theta}\bm{n})_{\partial M}=\sum_{e\in\mathcal{F}_{pr}^{\frac{1}{2}}}(\left\llbracket\bm{v}\right\rrbracket,\underline{I}_{h}^{e}\underline{\theta}\bm{n})_{e}=\|\bm{v}\|_{L^{2}(\Omega)}^{2}.\end{split}$ (4.38)
In addition, it follows from the triangle inequality, the boundedness of $\underline{\Pi}_{h}$ and the interpolation error estimates of $\underline{I}_{h}^{e}$ that
$\|\underline{\Pi}_{h}\underline{I}_{h}^{e}\underline{\theta}\|_{L^{2}(\Omega)}\leq\|\underline{\Pi}_{h}\underline{I}_{h}^{e}\underline{\theta}-\underline{I}_{h}^{e}\underline{\theta}\|_{L^{2}(\Omega)}+\|\underline{I}_{h}^{e}\underline{\theta}\|_{L^{2}(\Omega)}\leq C\|\underline{I}_{h}^{e}\underline{\theta}\|_{L^{2}(\Omega)}\leq C\|\underline{\theta}\|_{H^{1}(\Omega)}.$
On the other hand, we have from the definition of $\underline{\Pi}_{h}$, integration by parts and the trace inequality that
$\|\mathrm{div}\underline{\Pi}_{h}\underline{I}_{h}^{e}\underline{\theta}\|_{L^{2}(\Omega)}^{2}$ $=\sum_{M\in\mathcal{T}_{M}}(\mathrm{div}\Pi_{h}\underline{I}_{h}^{e}\underline{\theta},\mathrm{div}\underline{\Pi}_{h}\underline{I}_{h}^{e}\underline{\theta})_{M}=\sum_{M\in\mathcal{T}_{M}}(\underline{I}_{h}^{e}\underline{\theta}\bm{n},\mathrm{div}\underline{\Pi}_{h}\underline{I}_{h}^{e}\underline{\theta})_{\partial M}$
$=\sum_{e\in\mathcal{F}_{pr}^{\frac{1}{2}}}(\underline{I}_{h}^{e}\underline{\theta}\bm{n},\left\llbracket\mathrm{div}\underline{\Pi}_{h}\underline{I}_{h}^{e}\underline{\theta}\right\rrbracket)_{e}=\sum_{e\in\mathcal{F}_{dl}}(\left\llbracket\underline{I}_{h}^{e}\underline{\theta}\bm{n}\right\rrbracket,\mathrm{div}\underline{\Pi}_{h}\underline{I}_{h}^{e}\underline{\theta})_{e}$
$\leq C\sum_{e\in\mathcal{F}_{dl}}h_{e}^{-1/2}\|\left\llbracket\underline{I}_{h}^{e}\underline{\theta}\bm{n}\right\rrbracket\|_{L^{2}(e)}h_{e}^{1/2}\|\mathrm{div}\underline{\Pi}_{h}\underline{I}_{h}^{e}\underline{\theta}\|_{L^{2}(e)}$
$\leq C\|\underline{I}_{h}^{e}\underline{\theta}\|_{\Sigma,h}\|\mathrm{div}\underline{\Pi}_{h}\underline{I}_{h}^{e}\underline{\theta}\|_{L^{2}(\Omega)},$
which yields
$\|\mathrm{div}\underline{\Pi}_{h}\underline{I}_{h}^{e}\underline{\theta}\|_{L^{2}(\Omega)}\leq C\|\underline{I}_{h}^{e}\underline{\theta}\|_{\Sigma,h}.$
Therefore, the following estimate holds in view of (4.37) and (4.36)
$\|\underline{\Pi}_{h}\underline{I}_{h}^{e}\underline{\theta}\|_{\mathrm{div}}\leq C\|\underline{I}_{h}^{e}\underline{\theta}\|_{\Sigma,h}\leq C\|\bm{v}\|_{L^{2}(\Omega)}.$
This and (4.38) complete the proof by taking $\underline{w}=\underline{I}_{h}^{e}\underline{\theta}$.
@@ 4.2 Quasi-optimal approximation
Theorem 4.3.
Let $(\underline{\sigma},\bm{u},\gamma)$ be the exact solution and let $(\underline{\sigma}_{h},\bm{u}_{h},\gamma_{h})\in\underline{\Sigma}_{h}\times\bm{U}_{h}\times\Gamma_{h}$ be the discrete solution of (4.9)-(4.10) obtained by either Method 1 or Method 2. Assume that $(\underline{\sigma},\gamma)\in\underline{H}^{1}(\Omega)\times H^{1}(\Omega)$, then the following convergence error estimate holds
$\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}\leq C\Big{(}h|\underline{\sigma}|_{H^{1}(\Omega)}+h|\gamma|_{H^{1}(\Omega)}+\Big{(}\sum_{M\in\mathcal{T}_{M}}h_{M}^{2}\|\bm{f}\|_{L^{2}(M)}^{2}\Big{)}^{\frac{1}{2}}\Big{)}.$

### v2

@@ 4.1 Some elementary properties
$\sum_{e\in\mathcal{F}_{pr}^{\frac{1}{2}}}(\underline{w}_{h}\bm{n},\left\llbracket\bm{v}_{h}\right\rrbracket)_{e}+(\text{as}(\underline{w}_{h}),\eta_{h})=\sum_{e\in\mathcal{F}_{pr}^{\frac{1}{2}}}h_{e}^{-1}\|\left\llbracket\bm{v}_{h}\right\rrbracket\|_{L^{2}(e)}^{2}+\|\eta_{h}\|_{L^{2}(\Omega)}^{2},$
where we use $\mathrm{curl}(\bm{\rho}_{h})\bm{n}_{|e}=\bm{0}$ for any $e\in\mathcal{F}_{pr}^{\frac{1}{2}}$ if $\bm{\rho}_{h}=\bm{0}$ on $\partial M$. This, (4.15) and (4.17) complete the proof.
Lemma 4.3.
(inf-sup condition). For Method 2, there exists a positive constant $C$ such that
$\sup_{\bm{0}\neq\underline{w}\in\underline{\Sigma}_{h}}\frac{(\bm{v},\mathrm{div}\underline{\Pi}_{h}\underline{w})+(\textnormal{as}(\underline{w}),\eta)}{\|\underline{w}\|_{L^{2}(\Omega)}}\geq C(\|\bm{v}\|_{L^{2}(\Omega)}+\|\eta\|_{L^{2}(\Omega)})\quad\forall(\bm{v},\eta)\in\bm{U}_{h}\times\Gamma_{h}.$
Proof.
$\|\underline{I}_{h}^{e}\underline{\theta}\|_{L^{2}(\Omega)}\leq C\|\underline{\theta}\|_{H^{1}(\Omega)}.$ (4.19)
We can infer from (4.18) and (4.2) that
Then we have
$\begin{split}(\bm{v},\mathrm{div}\underline{\Pi}_{h}\underline{I}_{h}^{e}\underline{\theta})&=\sum_{M\in\mathcal{T}_{M}}(\bm{v},\underline{\Pi}_{h}\underline{I}_{h}^{e}\underline{\theta}\bm{n})_{\partial M}\\ &=\sum_{M\in\mathcal{T}_{M}}(\bm{v},\underline{I}_{h}^{e}\underline{\theta}\bm{n})_{\partial M}=\sum_{e\in\mathcal{F}_{pr}^{\frac{1}{2}}}(\left\llbracket\bm{v}\right\rrbracket,\underline{I}_{h}^{e}\underline{\theta}\bm{n})_{e}=\|\bm{v}\|_{L^{2}(\Omega)}^{2}.\end{split}$ (4.20)
We define $\underline{w}_{2}=\mathrm{curl}(\bm{\rho}_{h})$ with $g=\eta-\underline{I}_{h}^{e}\underline{\theta}$ (cf. Lemma 4.1). Then we have $\mathrm{div}\underline{w}_{2}=\bm{0}$ and an application of Lemma 4.1 yields
$(\text{as}(\underline{I}_{h}^{e}\underline{\theta}+\underline{w}_{2}),\eta)=\|\eta\|_{L^{2}(\Omega)}^{2}\quad\mbox{and}\quad\|\underline{w}_{2}\|_{L^{2}(\Omega)}\leq C\|\bm{v}\|_{L^{2}(\Omega)}.$
This and (4.20) complete the proof by taking $\underline{w}=\underline{I}_{h}^{e}\underline{\theta}+\underline{w}_{2}$.
@@
We remark that the proof for Lemma 4.2 and Lemma 4.3 can be easily extended to proving the inf-sup condition for Method 1. The only difference is to define the Stokes pair used in Lemma 4.1 for each $D,D\in\mathcal{T}_{D}$; indeed, we treat $D$ as the “macro-element” in this case.
@@
(existence and uniqueness). There exists a unique solution to (3.2)-(3.3). In addition, there exists a positive constant $C$ independent of the mesh size such that the following estimate holds
@@
Since (3.2)-(3.3) is a square linear system, the uniqueness of solution implies the existence. Setting $\bm{f}=\bm{0}$ in (4.21) implies $\underline{\sigma}_{h}=\underline{0},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{u}_{h}=\bm{0}}$ and $\gamma_{h}=0$.
@@
4.2 The convergence error analysis
Lemma 4.6.
Let $(\bm{u},p)$ be the exact solution and let $(\bm{u}_{h},p_{h})\in\bm{U}_{h}\times P_{h}$ be the discrete solution of (4.6)-(4.7). Then, the following error equations hold
$(\mathcal{A}(\underline{\sigma}-\underline{\sigma}_{h}),\underline{w})+(\bm{u}-\bm{u}_{h},\mathrm{div}\underline{\Pi}_{h}\underline{w})+(\gamma-\gamma_{h},\textnormal{as}(\underline{w}))$ $=(\mathcal{A}\underline{\sigma}+\gamma\underline{\delta},\underline{w}-\underline{\Pi}_{h}\underline{w})\quad\forall\underline{w}\in\underline{\Sigma}_{h},$ (4.23)
$(\mathrm{div}\underline{\Pi}_{h}(\underline{\sigma}-\underline{\sigma}_{h}),\bm{v})$ $=0\quad\forall\bm{v}\in P_{h},$ (4.24)
$(\textnormal{as}(\underline{\sigma}-\underline{\sigma}_{h}),\xi)$ $=0\quad\forall\xi\in\Gamma_{h}.$ (4.25)